A company desires to sell a sufficient quantity of products to earn a profit of $260000. If the unit sales price is $20, unit variable cost is $12, and total fixed costs are $800000, how many units must be sold to earn net income of $260000?

A. 72500 units

B. 113000 units

C. 216250 units

D. 132500 units
Answer #1
Desired income: $ 260,000
Fixed cost: $ 800,000
Desired contribution :Desired income+Fixed cost =260,000+800,000 =1060,000
Selling price per unit =20 per unit
Variable cost per unit: 12 per unit
Contribution margin per unit =Selling price -Variable cost =20-12 = $ 8 per unit
Target sales in units: Desired contribution/ Contribution margin per unit
1060,000/8 = 132,500 units